# Elastic properties of three-dimensional Yukawa or dust crystals from molecular dynamic simulations

### v1

@@ IV Summary
I have calculated bulk modulus, shear modulus, and Poisson’s ratio of a cubic Yukawa or dusty crystal in the molecular dynamic simulations. These properties are computed by deforming (compressing/expanding) the dust crystal along different directions. The characteristics of elastic properties as a function of shielding parameter and strong coupling parameter have been studied. The shear modulus has also been calculated along the crystallographic directions of 100 and 110, and it is found that the values are different at small screening parameters. Furthermore, I have calculated sound and shear wave velocities using bulk and shear modulus and compared them with the velocities obtained from the slope of longitudinal and transverse wave dispersion relations. It is found that the sound and shear velocities obtained using bulk and shear modulus agree with the previously observed experimental values.
Strongly coupled dusty plasmas exhibit solid-like traits, but they are extremely soft. As a result, the bulk modulus and shear modulus of solid dusty plasmas are much smaller than those of typical solids, such as metals. In fact, the bulk modulus is much larger compared to the shear modulus, which effectively renders dusty plasma as incompressible. This is because it is easier to change shape than to compress a dusty plasma system.
@@ Acknowledgements.
The author thanks Dr. Attila Cangi (CASUS-HZDR, Görlitz, Germany) and Dr. Svetoslav Nikolov (SNL, Albuquerque, USA) for helpful discussions. The computations were performed on a Bull Cluster at the Center for Information Services and High Performance Computing (ZIH) at Technische Universität Dresden, on the cluster Hemera at Helmholtz-Zentrum Dresden-Rossendorf (HZDR).

### v2

@@ IV Summary
I have calculated bulk modulus, shear modulus, and Poisson’s ratio of a cubic Yukawa or dusty-crystal using molecular dynamic simulations. These properties have been computed by deforming (compressing/expanding) the dust crystal along different directions. The characteristics of elastic properties as a function of shielding parameter and strong coupling parameter have been presented. The shear modulus of the Yukawa system has also been determined along the crystallographic directions of 100 and 110. Furthermore, I have obtained sound and shear wave velocities using bulk and shear modulus and compared them with the velocities obtained from the slope of longitudinal and transverse wave dispersion relations. It has been found that the sound and shear velocities obtained using bulk and shear modulus agree with the experimental values. It has also been found from my computations that the bulk modulus is much larger than the shear modulus, which effectively renders dusty plasma as incompressible. This is because it is easier to change shape than to compress a dusty plasma system. I think that my findings on elastic properties can also apply to other systems, such as charged colloids, electrolytes, and strongly coupled electron-ion plasmas.
@@ Acknowledgements.
The author thanks Dr. Attila Cangi (CASUS, Görlitz, Germany) and Dr. Svetoslav Nikolov (SNL, Albuquerque, USA) for helpful discussions. The computations were performed on a Bull Cluster at the Center for Information Services and High Performance Computing (ZIH) at Technische Universität Dresden, on the cluster Hemera at Helmholtz-Zentrum Dresden-Rossendorf (HZDR), and on TACC Frontera at University of Texas at Austin.
Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.